A flywheel of an engine has a mass of 1250 kg and radius of gyration 600 mm. Find the angular acceleration of the wheel, when it is subjected to a torque of 12 500 N-m.


Ans. 27.75 rad/s)


1
Expert's answer
2021-02-18T18:40:36-0500
"\tau=I\alpha,""\tau=mk^2\alpha,""\alpha=\dfrac{\tau}{mk^2}=\dfrac{12500\ N\cdot m}{1250\ kg\cdot(0.6\ m)^2}=27.77\ \dfrac{rad}{s^2}."